OpenStudy (anonymous):

-4x + 2y = 8 4x – 3y = -10 i am very confused on what to do i have some of this worked but i am stuck

OpenStudy (anonymous):

do you have to find both the x and y values?

OpenStudy (anonymous):

What methods have you learned so far to solve systems of equations?

OpenStudy (anonymous):

yes do you want me to show you what i already have worked?

OpenStudy (anonymous):

yes please it will help a lot.

OpenStudy (anonymous):

-4x+2y=8 4x-3y=-10 -4x+4x=0 2y-3y=-y 8-10=-2 -y=-2 y=2 and then i dont know what else to do or how to check my work

OpenStudy (anonymous):

now plug in y=2 to one of the two equations and solve for x. For example -4x+2y=8 , y=2 -4x+2(2)=8 -4x+4=8 Can you do the rest?

OpenStudy (anonymous):

Your solution will be x=? , y=2

OpenStudy (anonymous):

x=? is whatever you find for x

OpenStudy (anonymous):

x=-1 and positive 4

OpenStudy (anonymous):

do you know how to do a check step?

OpenStudy (anonymous):

-4x + 2y=8 answer =y/2 -2

OpenStudy (anonymous):

4x – 3y = -10 answer = x =3y/4 -5/2